OpenStudy (anonymous):

The half-life of carbon-14 is approximately 6000 years. Determine how much of 100 kg of this substance will remain after 12,000 years.

OpenStudy (anonymous):

@terenzreignz

OpenStudy (anonymous):

@amistre64 @SolomonZelman

OpenStudy (solomonzelman):

Dang I forgot how to do these. -:( I am replying to get a notification and to later view this question. (You guys are probably thinking that I am typing the greatest tip/solution in the world right now... well , not really)

terenzreignz (terenzreignz):

I was... going to ask you if you were up to the task, Zelman XD Okay, listen up... What do you know about half life, @hello1213 ?

OpenStudy (anonymous):

I know the formula for half-life decay... it is\[N=N _{0}(\frac{ 1 }{ 2 })^\frac{ t }{ h }\]

terenzreignz (terenzreignz):

Well, It's just a matter of plugging in :D t here is the time that passed and h is the half life \(N_0\) is the amount of the substance at the start. Plug in and be done ^_^

OpenStudy (anonymous):

Well I think that's where I am having trouble not plugging in, but simply solving the equation.

terenzreignz (terenzreignz):

Okay... here we go... \[\Large N = N_0 \left(\frac12\right)^{\frac t h}\] Let's plug in, one at a time. What is \(N_0 = \color{red}?\)

OpenStudy (anonymous):

100?

terenzreignz (terenzreignz):

Right. \[\Large N =\color{blue}{100} \left(\frac12\right)^{\frac t h}\] What is \(t =\color{red}?\) Or, in other words... how much time has passed?

OpenStudy (anonymous):

12,000

terenzreignz (terenzreignz):

Correct again :D \[\Large N =100\left(\frac12\right)^{\frac {\color{blue} {12000}} h}\] What about h, the half life?

OpenStudy (anonymous):

so h would be 6000.

terenzreignz (terenzreignz):

Okay, we finally have N in terms of numbers only... lol \[\Large N = 100 \left(\frac12\right)^{\frac {12000}{6000}}\] But... first things first, what is \(\Large \frac{12000}{6000}\) equal to?

OpenStudy (anonymous):

2

terenzreignz (terenzreignz):

Right... \[\Large N = 100 \left(\frac12\right)^2\] Can you take it from here?

OpenStudy (anonymous):

Ohhh, I didn't realize 2 was a power before, so I was writing it wrong, and multiplying 2 by 1/2, when 2 was a power... :/... Yes so the answer is N=25!

terenzreignz (terenzreignz):

That is correct. Well done ^_^

OpenStudy (anonymous):

Thanks again... You always make these seemingly difficult problems easy!

terenzreignz (terenzreignz):

I do try ^_^